Near the equator, The field lines of the Earth's magnetic field are approximately horizontal to the north and the magnetic flux density is 50 μT. How large and in what direction must the speed of the electron be in order for the force caused by the magnetic field to cancel out the gravity acting on the electron?




1
Expert's answer
2021-08-31T08:58:57-0400

"F_l=F,"

"F_l=evB,"

"F=mg,"

"evB=mg,"

"v=\frac{mg}{eB}=1.1\cdot 10^{6}~\frac ms."